Fermi National Laboratory
 

Accelerator Update for June 27, 2006 - June 28, 2006

 

Tuesday June 27

The day shift began with Operations monitoring store 4799 and a Recycler stash, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, with the Antiproton Source (Pbar) stacking, with the Linac, Booster, Main Injector (MI), Switchyard (SY), and Recycler operating normally, and with MiniBooNE and NuMI taking beam.

At 8:19 AM, the NuMI Horn power supply tripped off and then at 8:27 AM there was a target and baffle trip. Both of these caused the NuMI beam permit to trip. Operations contacted an expert who directed that these trips be reset.

At 8:34 AM, a Linac RF station (KRF2) tripped off on a LCW no flow indication. Operators reset the station.

At 9:59 AM, the Neutron Therapy Facility (NTF) reported water in the lower Linac gallery. Operators investigated and then called the Duty Mechanic. He found a pinhole leak in a chilled water pump discharge pipe. He switched to a different pump.

Operations established beam to the Meson MTest experiment at 10:40 AM.

At 11:08 AM, A Linac expert reported that KRF2 wasn’t regulating its cavity phase properly. Beam was intermittent while they made repairs.

At 11:41 AM, the MTest beam permit tripped off with no status. Operations reset the permit and then a power supply tripped off (HP3DS). It reset okay.

At 1:35 PM, the MTest permit tripped due to a power supply undercurrent indication. It reset okay.

At 1:43 PM, the MiniBooNE beam permit tripped off due to a power supply undercurrent indication. It reset okay.

At 1:47 PM, Operations reported that a TeV sector C4 air compressor had tripped off earlier. It reset okay.

Operation began preparing the Recycler at 2:10 PM to stash antiprotons.

Pbar experts began a one hour 8 GeV beamline lattice study at 2:44 PM.

The stash ended at 3:59 PM.

Operations terminated store 4799 at 4:01 PM.

Operations made CDF and D0 ready for an access at 4:12 PM.

Experts opened the TeV VCBs at 4:32 PM to investigate a sector A2 dump LCW leak.

MTest resumed taking beam around 4:31 PM.

A TeV expert reported at 6:06 PM that he changed out a leaky shunt module on the A2 dump. Operations performed a dump test at 7:39 PM.

At 8:11 PM, operators found that a Meson front-end power supply had failed at EAD. This shut off beam to MTest. Operators contacted experts.

Operations put the TeV through a dry squeeze at 8:25 PM.

Operations put the TeV into shot setup at 8:57 PM.

MTest gave up for the night at 10:05 PM.

An expert replaced a CAMAC crate for Meson at 11:29 PM.

Operations established store 4802 at 11:57 PM with an initial luminosity of 98.85E30.

Wednesday June 28

The midnight shift began with Operations monitoring store 4802 and the stash, with Pbar stacking, with normal Linac, MI, Recycler, and Booster operations, and with MiniBooNE and NuMI taking beam.

At 12:50 AM, a MiniBooNE expert reported numerous Horn trips. Operators investigated and found a vacuum pump tripped off (B:IPQ16). It reset okay.

The Recycler began preparing to stash at 1:25 AM.

Operations reported at 1:34 AM that the MiniBooNE trips were due to a timeline problem.

At 2:17 AM, the Booster IPQ16 vacuum pump tripped off for the first of three times in eighteen minutes. It reset okay each time.

The stash ended and Pbar resumed stacking at 2:21 AM.

The Recycler began preparing to stash at 6:27 AM. The stash ended and Pbar resumed stacking at 7:21 AM.

The Plans for Wednesday

The plans for today are to stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/28/2006    email Fermilab